# Polarization Signatures of Quasi-Periodic Oscillations in Simulated Tilted, Truncated Disks

### v2

@@ 3.1.1 Spectral Components
One of the powerful things we can do with a ray-tracing code like Pandurata is track each photon packet’s history from creation in the photosphere to its collection at the camera. We can then use this information to break the flux and polarization spectra down according to the history of the different photon packets, as we do in Figures 4 and 5. In this paper, we track “Disk” photons that originate in the photosphere and escape to the observer without scattering, photons that scatter once off the disk (“Disk Scatter”, also called returning radiation), photons that scatter once in the corona (“Corona Scatter”), and photons that scatter once off the disk and once in the corona (“Disk & Corona”). When working with the Stokes parameters (see equations 2.2), we can simply add the frequency-dependent $I_{\nu}$, $Q_{\nu}$, and $U_{\nu}$ from all the relevant photon packets to get the total spectrum and polarization signal, as given in Figure 2 and depicted in Figure 4 as solid black lines.
Figure 4: Flux (top panel), polarization degree (middle panel), and polarization angle (bottom panel) as a function of energy for a nearly face-on observer ($i=18^{\circ}$) broken down by the history of the photon packets. These data are for the untilted, truncated disk simulation (a9b0L4) at $t=24,000\,t_{g}$. “Disk” photons are created at the photosphere, “Disk Scatter” photons scatter once off the disk, “Corona Scatter” photons scatter once in the corona, and “Disk & Corona” photons scatter once each off the disk and corona. The solid black line in each panel gives the total value when all the components are properly weighted.
@@
In Figure 4 (the nearly face-on case), we see that the disk contribution dominates the flux spectrum (top panel) to slightly past the thermal cutoff ($\sim 5$ keV), beyond which the corona scatter and disk & corona components contribute roughly equally to the power-law tail. Not surprisingly then, in the polarization measures, the disk component factors most heavily up to that same cutoff, beyond which, the corona scatter photons begin to have a greater influence on the polarization. (The disk & corona component always has a very small polarization degree, close to 0.)
For the nearly edge-on case in Figure 5, the situation is very different. The corona scatter photons dominate the flux spectrum over the entire energy range. Thus, they also largely dominate the polarization measures, although the disk photons keep the polarization from being as high as it might otherwise be at low energies ($\lesssim 1$ keV). In both cases we see that the disk scatter component is characterized by a nearly vertical polarization angle (±$\pm 90^{\circ}$), due to the large-angle scattering experience by returning radiation (Schnittman & Krolik, 2009).
@@ 3.1.2 Time Dependence
Lastly, we look at how the flux and polarization measures vary as a function of time for the untilted case (Figure 6). We find that, after some initial period of adjustment lasting $\lesssim 20,000\,t_{g}$, all of the measures reach nearly steady values (dependent only on inclination). This is consistent with the lack of time variability seen in this simulation’s hydrodynamic variables (Bollimpalli et al., 2023, 2024). The different sensitivity ranges of IXPE (2-8 keV) and XPoSat (8-30 keV) explain why the curves look different for those two instruments even at the same inclination.
Figure 6: Flux (top panel), polarization degree (middle panel), and polarization angle (bottom panel) plotted as a function of time for three different observer inclinations integrated over the IXPE sensitivity range and one inclination for XPoSat for our untilted simulation (a9b0L4). After a transient period ($\lesssim 20,000\,t_{g}$), all of the measures settle into roughly constant values.
@@ 3.2 Tilted, Truncated Disk
Our ultimate goal in this work is to demonstrate that QPOs are observable through their time-varying polarization signatures. We consider both a low-frequency QPO attributed to the precession of a tilted, thick-disk component and higher frequency QPOs that show up spontaneously in our tilted simulations. For our tilted, truncated disk simulation (a9b15L4), we focus on the part of the simulation that covers the first quarter of a precession period (about $25,000\,t_{g}$). Unfortunately, it is difficult in this limited period to discern the effects of precession in the polarization light curves, owing to confusion with the initial transient period of the simulation, which lasts almost $20,000\,t_{g}$, as seen for the untilted case (Figure 6).
Therefore, rather than looking for the time variability due to precession in the light curves of a fixed observer, we instead focus on observations at a single time, but from different azimuths (moving the camera to simulate precession). In order to properly capture the effect of the inner region precessing, it is important to keep the observer inclination fixed with respect to the symmetry axis of the outer disk (as nearly as possible given our finite binning). In this section, we keep that inclination at $i=72^{\circ}$. Note that this means we are changing the inclination of the observer with respect to the black hole spin axis and symmetry axis of the inner torus, which is why we expect the polarization signature to vary. In the next section (3.3), we present results of a simulation where we follow an isolated torus for more than a full precession cycle and can therefore focus on a light curve from a single observer azimuth.
Before considering the azimuthal dependence of the polarization measures, we first comment on the overall flux from our tilted, truncated disk (Figure 7, top panel). Comparing with the $i=72^{\circ}$ curve in Figure 2, we see that the peak flux in the tilted case is higher for most azimuths, and also that the spectrum is harder (peaks at higher energies). This may be partly due to extra dissipation associated with tilted disks (Fragile & Blaes, 2008; Generozov et al., 2014), but is mostly attributed to the much greater coronal power in the tilted case (compare the top and bottom panels of Figure 1).
@@
More interestingly, there is a clear azimuthal dependence to the flux, especially above 1 keV, indicating that this QPO could be detected in both the IXPE and XPoSat bands. The polarization signatures (middle and bottom panels of Figure 7) also vary greatly with azimuth. The polarization degree appears about equally variable in both the IXPE and XPoSat bands, whereas the polarization angle is definitely more variable in the XPoSat band, swinging from about $50^{\circ}$ for the $\phi=315-360^{\circ}$ bin to $-70^{\circ}$ for the $45-90^{\circ}$ bin. The polarization angle can even have different signs for the same azimuth depending on the energy band (e.g., the blue and orange curves). Similar energy dependence of the polarization angle has been observed for at least one active galactic nucleus source (Gianolli et al., 2024), although entirely within the IXPE band and with the polarization degree rising, rather than falling, with energy. These results demonstrate the potential of polarization to diagnose tilted disks and precession. We would expect the variability over one precession cycle to be similar to that exhibited in Figure 7, which we confirm for an isolated, tilted torus in Section 3.3.
@@ 3.2.1 Spectral Components
To better understand the results of Figure 7, we can break down the observations into their spectral components as we did in Figure 4. We focus, in Figure 8, on the $\phi=45-90^{\circ}$ azimuthal bin. We find that the disk component almost matches the corona scattering up to almost 10 keV, beyond which the corona scatter and disk & corona components dominate, somewhat similar to Figure 4. Accordingly, these same three components control the polarization measures. Here, however, is where the differences from the untilted case become most apparent. The degree and angle of polarization show interesting variations with energy right through the sensitivity ranges of IXPE and XPoSat. These variations are driven mainly by changes in the corona scatter contribution, which is the component that should be most sensitive to tilt.
@@ 3.2.2 High-frequency QPO
One of the most remarkable results of our tilted, truncated disk simulations is that, along with the precession of the inner torus, they also spontaneously develop high-frequency QPOs (Bollimpalli et al., 2024). In terms of hydrodynamic variables, these QPOs are most apparent in the vertical velocity component of the gas, as it seems the presence of tilt excites vertical oscillations of the inner torus. For the a9b15L4 simulation, this QPO has a period of approximately $500\,t_{g}$ (approximately 40 Hz for a $10M_{\odot}$ black hole), which can be seen in Figure 9 (middle panel) where we plot the time evolution of the disk’s mid-plane angle, $\theta_{\mathrm{mid}}$, measured at a fixed azimuthal angle near the inner edge of the disk. The period of oscillations in that curve closely match the period of oscillations seen in the polarization light curves covering the same simulation time, especially in the higher-energy XPoSat band.
Figure 9: Same as Fig. 6, except for a short segment of our tilted, truncated disk simulation (a9b15L4) in the $315^{\circ}<\phi\leq 360^{\circ}$ bin. The middle panel also includes the time evolution of the disk’s mid-plane angle, $\theta_{\mathrm{mid}}$, measured at a fixed azimuth, $\phi=0$, for $r=5\,r_{g}$ (solid, black), showing the vertical oscillations of the inner torus.
@@ 3.3 Isolated, Tilted Torus
The last simulation we consider is of an isolated, tilted torus. We include this case because it is the only one of our simulations that evolved for a full precession period. Thus, we can create a complete map between the polarization measures and precession phase. This is done in Figure 10, where, like Figure 6, we plot the flux, polarization degree, and polarization angle as a function of time. However, unlike Figure 6 where the lines are basically flat after some initial period of adjustment, we see prominent fluctuations in the polarization measures. The polarization angle, in particular, clearly traces the two precession periods of the simulation for all observer inclinations and energy ranges. The polarization degree is less well synchronized with the precession, though the intermediate inclination ($i=45^{\circ}$) does show roughly two periods of oscillation, though with evidence for a possible higher-frequency oscillation added on top of the precession.
Figure 10: Same as Fig. 6, except for the isolated, tilted torus and only including photon packets arriving in the $315^{\circ}<\phi\leq 360^{\circ}$ azimuthal bin. This plot covers roughly two full precession periods, which are easily seen in the polarization angle for all inclinations and energy bands.
Indeed, if we focus on a higher-time-resolution interval, as in Figure 11, then we clearly see another QPO, this one with a period of $\sim 200\,t_{g}$. This is interesting, as no high-frequency QPO was previously reported for this simulation. This may be because Bollimpalli et al. (2024) focused primarily on radial and vertical oscillation modes, whereas this QPO seems to be associated with the orbital modulation of an $m=1$, non-axisymmetric mode that arises from the Papaloizou-Pringle instability (PPI; Papaloizou & Pringle, 1984). This particular high-frequency QPO may not be a realistic mode for a full accretion disk, as the PPI applies most strongly to small, isolated tori, but that is not really the point. The point is that we have demonstrated that polarization measures are sensitive to QPOs with a variety of frequencies and modes.
@@ 4.1 Comparison to Previous Work
Broadly speaking, our results show similarities to previous studies (e.g., Schnittman & Krolik, 2009), especially in cases where the disk component dominates the flux (i.e., our untilted, nearly face-on results). In those cases, the spectrum is consistent with a multi-temperature blackbody that is well modeled by Novikov & Thorne (1973). The polarization degree is modest (only a few percent), and the polarization angle is close to zero (i.e., horizontal). When including the returning radiation, we expect a transition from horizontal to vertical polarization just beyond the thermal peak, as the large scattering angles (and thus high degree of vertical polarization) of the deflected disk photons begin to dominate the polarization signal. As described in Schnittman & Krolik (2009), this transition occurs at lower energies for lower inclinations, as the direct disk contribution is inherently less polarized, so easier to “overcome” by the scattered disk contribution.
@@ 4.2 Prospects for Detection
In this work, we have shown that various QPO mechanisms, including global precession of a tilted torus, orbital modulation by a non-axisymmetric mode, and vertical oscillations in a tilted, truncated disk, can all produce notable changes in polarization. Our hope is that these polarization modulations may be detectable with a dedicated X-ray polarimeter, such as IXPE (Weisskopf et al., 2022) or XPoSat (Paul, 2022). However, we must acknowledge that this would be a very challenging measurement to make in practice. The main difficulty has to do with timescales. Even the low-frequency, type-C QPO has frequencies in the range of 0.1-10 Hz, while polarization measurements typically require long exposure times (tens or hundreds of kiloseconds for IXPE). This is obviously many orders of magnitude longer than the QPO period.
@@ 5 Conclusion
We have post-processed the untilted and tilted, truncated black hole accretion disk and isolated, tilted torus simulations of Bollimpalli et al. (2024) using the polarization-capable, Monte-Carlo radiation transport code Pandurata  (Schnittman & Krolik, 2013). Our results show reasonable agreement with previous work in that, whenever the disk component dominates, we find a roughly thermal spectrum with an emission peak at a few keV, appropriate for a black hole with $a_{*}=0.9$, disk images that appear nearly symmetric, and polarization degrees that are generally low.
@@
For our untilted, truncated disk, we find very little azimuthal or time dependence, as expected, though we do find a strong inclination dependence. For a nearly edge-on observation, the spectrum can be quite hard, with coronal scattering dominating at all energies. The polarization degree can also be large ($\gtrsim 5$%) in such cases.
@@
We gratefully acknowledge support from SC NASA EPSCoR RGP 2024 and NASA under award No 80NSSC24K0900. DAB acknowledges support from IIT-Indore, through a Young Faculty Research Seed Grant (project: ‘INSIGHT’; IITI/YFRSG/2024-25/Phase-VII/02).

### v3

@@ 3.1.1 Spectral Components
One of the powerful things we can do with a ray-tracing code like Pandurata is track each photon packet’s history from creation in the photosphere to its collection at the camera. We can then use this information to break the flux and polarization spectra down according to the history of the different photon packets, as we do in Figures 4 and 5. In this paper, we track “Disk” photons that originate in the photosphere and escape to the observer without scattering, photons that scatter once off the disk (“Reflection”, also called returning radiation), photons that scatter only in the corona (“Corona Scatter”), and photons that scatter both off the disk and in the corona (“Reflection + Corona”). When working with the Stokes parameters (see equations 2.2), we can simply add the frequency-dependent $I_{\nu}$, $Q_{\nu}$, and $U_{\nu}$ from all the relevant photon packets to get the total spectrum and polarization signal, as given in Figure 2 and depicted in Figures 4 and 5 as solid black lines.
Figure 4: Flux (top panel), polarization degree (middle panel), and polarization angle (bottom panel) as a function of energy for a nearly face-on observer ($i=18^{\circ}$) broken down by the history of the photon packets. These data are for the untilted, truncated disk simulation (a9b0L4) at $t=24,000\,t_{g}$. “Disk” photons are created at the photosphere, “Reflection” photons scatter once off the disk, “Corona Scatter” photons scatter in the corona, and “Reflection + Corona” photons scatter off both the disk and corona. The solid black line in each panel gives the total value when all the components are properly weighted.
@@
In Figure 4 (the nearly face-on case), we see that the disk, corona scatter, and reflection + corona photons contribute roughly equally to the flux spectrum (top panel) up to the thermal cutoff ($\sim 1$ keV), beyond which the reflection + corona component creates a power-law tail. Not surprisingly then, in the polarization measures, the disk, corona scatter, and reflection + corona components each contribute up to that same cutoff, beyond which, the reflection + corona mostly dominates.
For the nearly edge-on case in Figure 5, the situation is somewhat different. The corona scatter and reflection + corona photons dominate the flux spectrum over the entire energy range. Thus, they also largely dominate the polarization measures, although the disk photons keep the polarization degree from being as high as it might otherwise at low energies ($\lesssim 1$ keV). In both cases we see that the reflection component is characterized by a nearly vertical polarization angle ($90^{\circ}$), due to the large-angle scattering experienced by returning radiation (Schnittman & Krolik, 2009).
@@ 3.1.2 Time Dependence
Lastly, we look at how the flux and polarization measures vary as a function of time for the untilted case (Figure 6). We find that, after some initial period of adjustment lasting $\lesssim 15,000\,t_{g}$, all of the measures reach nearly steady values (dependent only on inclination). This is consistent with the lack of time variability seen in this simulation’s hydrodynamic variables (Bollimpalli et al., 2023, 2024). The different sensitivity ranges of IXPE (2-8 keV) and XPoSat (8-30 keV) explain why the curves look different for those two instruments even at the same inclination.
Figure 6: Flux (top panel), polarization degree (middle panel), and polarization angle (bottom panel) plotted as a function of time for three different observer inclinations integrated over the IXPE sensitivity range and one inclination for XPoSat for our untilted simulation (a9b0L4). After a transient period ($\lesssim 15,000\,t_{g}$), all of the measures settle into roughly constant values.
@@ 3.2 Tilted, Truncated Disk
Our ultimate goal in this work is to demonstrate that QPOs are observable through their time-varying polarization signatures. We consider both a low-frequency QPO attributed to the precession of a tilted, thick-disk component and higher frequency QPOs that show up spontaneously in our tilted simulations. For our tilted, truncated disk simulation (a9b15L4), we focus on the part of the simulation that covers the first quarter of a precession period (about $25,000\,t_{g}$). Unfortunately, it is difficult in this limited period to discern the effects of precession in the polarization light curves, owing to confusion with the initial transient period of the simulation, which lasts almost $15,000\,t_{g}$, as seen for the untilted case (Figure 6).
Therefore, rather than looking for the time variability due to precession in the light curves of a fixed observer, we instead focus on observations at a single time, but from different azimuths (moving the camera to simulate precession). In order to properly capture the effect of the inner region precessing, it is important to keep the observer inclination fixed with respect to the symmetry axis of the outer disk (as nearly as possible given our finite binning) [^4] We also tried keeping the observer inclination fixed with respect to the black hole spin axis, but this resulted in much larger variability than what is represented in Figure 7 due to what is in effect precession of both the inner torus and outer disk.. In this section, we keep that inclination at $i=72^{\circ}$. Note that this means we are changing the inclination of the observer with respect to the black hole spin axis and symmetry axis of the inner torus, which is why we expect the polarization signature to vary. In the next section (3.3), we present results of a simulation where we follow an isolated torus for more than a full precession cycle and can therefore focus on a light curve from a single observer azimuth.
Before considering the azimuthal dependence of the polarization measures, we first comment on the overall flux from our tilted, truncated disk (Figure 7, top panel). Comparing with the $i=72^{\circ}$ curve in Figure 2, we see that the peak flux in the tilted case is higher for all azimuths, and also that the spectrum is much harder (peaks at higher energies). This may be partly due to extra dissipation associated with tilted disks (Fragile & Blaes, 2008; Generozov et al., 2014), but is mostly attributed to the much greater coronal power in the tilted case (compare the top and bottom panels of Figure 1).
@@
More interestingly, there is a clear azimuthal dependence to the flux, especially above 1 keV, indicating that this QPO could be detected in both the IXPE and XPoSat bands. The polarization signatures (middle and bottom panels of Figure 7) also vary greatly with azimuth. The polarization degree is more variable in the XPoSat band, while the polarization angle is extremely variable in both bands, swinging from about $60^{\circ}$ for the $\phi=135-180^{\circ}$ bin to $100^{\circ}$ for the $45-90^{\circ}$ bin. Similar energy dependence of the polarization angle has been observed for at least one active galactic nucleus source (Gianolli et al., 2024). These results demonstrate the potential of polarization to diagnose tilted disks and precession. We would expect the variability over one precession cycle to be similar to that exhibited in Figure 7, which we confirm for an isolated, tilted torus in Section 3.3.
@@ 3.2.1 Spectral Components
To better understand the results of Figure 7, we can break down the observations into their spectral components as we did in Figures 4 and 5. We focus, in Figure 8, on the $\phi=45-90^{\circ}$ azimuthal bin. We find that the spectrum is dominated by corona scatter and reflection + corona scatter at all energies, similar to Figure 5. However, unlike that case, the spectral components here extend well beyond 10 keV. Because of the prominence of these two components, they naturally control the polarization measures across the whole spectrum. The polarization degree shows similar variation with energy through the IXPE and XPoSat sensitivity ranges to our untilted simulation, while the polarization angle has shifted to a notably larger value ($\sim 110^{\circ}$ vs. $\sim 60^{\circ}$), owing to the precession of the inner disk and the greater scattering experienced by the tilted disk.
@@ 3.2.2 High-frequency QPO
One of the most remarkable results of our tilted, truncated disk simulations is that, along with the precession of the inner torus, they also spontaneously develop high-frequency QPOs (Bollimpalli et al., 2024). In terms of hydrodynamic variables, these QPOs are most apparent in the vertical velocity component of the gas, as it seems the presence of tilt excites vertical oscillations of the inner torus. For the a9b15L4 simulation, this QPO has a period of approximately $500\,t_{g}$ (or frequency of approximately 40 Hz for a $10M_{\odot}$ black hole), which can be seen in Figure 9 (middle panel) where we plot the time evolution of the disk’s mid-plane angle, $\theta_{\mathrm{mid}}$, measured at a fixed azimuthal angle near the inner edge of the disk. Similar period oscillations are seen in the polarization light curves covering the same simulation time (Figure 9 middle and bottom panels). This is strong confirmation that polarization can, in principle, be used to detect and measure QPOs.
Figure 9: Same as Fig. 6, except for a short segment of our tilted, truncated disk simulation (a9b15L4) in the $315^{\circ}<\phi\leq 360^{\circ}$ bin. The middle panel also includes the time evolution of the disk’s mid-plane angle, $\theta_{\mathrm{mid}}$, measured at a fixed azimuth ($\phi=0^{\circ}$) for $r=5\,r_{g}$ (solid, black), showing the vertical oscillations of the inner torus. Note that we have left off the XPoSat light curve for clarity.
@@ 3.3 Isolated, Tilted Torus
The last simulation we consider is of an isolated, tilted torus. We include this case because it is the only one of our simulations that has evolved for a full precession period ($\sim 4,500\,t_{g}$) [^5] The isolated torus is able to precess much faster than the truncated disk torus in simulation a9b15L4 because accretion between the two components in the latter case significantly retards precession (see Bollimpalli et al., 2023, for more details). Thus, we can create a complete map between the polarization measures and precession phase. Thus, unlike Section 3.2 where we used a single simulation snapshot and moved the observer around in azimuth, in this section, we keep the observer fixed at a single azimuth. This is done in Figure 10, where, like Figure 6, we plot the flux, polarization degree, and polarization angle as a function of time. However, unlike Figure 6 where the lines are basically flat after some initial period of adjustment, in Figure 10 we see prominent fluctuations in the flux and polarization measures. The polarization angle, in particular, clearly traces the first precession period of the simulation for all observer inclinations. The polarization degree and the late-time polarization angle are less well synchronized with the precession, though the intermediate inclination ($i=45^{\circ}$) does show roughly two periods of oscillation. Together with Section 3.2, these results confirm that polarization will be sensitive to the precession of a disk component.
Figure 10: Same as Fig. 9, except for the isolated, tilted torus. This plot covers about 1.7 precession periods, which are most visible in the flux and polarization angle.
Figure 10 also shows evidence for a possible higher-frequency oscillation added on top of the precession. Indeed, if we focus on a higher-time-resolution interval, as in Figure 11, then we clearly see another QPO, this one with a period of $\sim 200\,t_{g}$. This is interesting, as no high-frequency QPO was previously reported for this simulation. This may be because Bollimpalli et al. (2024) focused primarily on radial and vertical oscillation modes, whereas this QPO seems to be associated with the orbital modulation of an $m=1$, non-axisymmetric mode that may arise from the Papaloizou-Pringle instability (PPI; Papaloizou & Pringle, 1984). This particular high-frequency QPO may not be a realistic mode for a full accretion disk, as the PPI applies most strongly to small, isolated tori. However, that is not really the point; the point is that we have demonstrated that polarization measures are sensitive to QPOs with a variety of frequencies and modes.
@@ 4.1 Comparison to Previous Work
Broadly speaking, our results show similarities to previous semi-analytic studies (e.g., Schnittman & Krolik, 2009), especially in cases where the disk component dominates the flux (i.e., our untilted, nearly face-on results). In those cases, the spectrum is consistent with a multi-temperature blackbody that is well modeled by Novikov & Thorne (1973). The polarization degree is modest (only a few percent), and the polarization angle is close to zero (or $180^{\circ}$, i.e., horizontal). When including the returning radiation, we expect a transition from horizontal to vertical polarization just beyond the thermal peak, as the large scattering angles (and thus high degree of vertical polarization) of the deflected disk photons begin to dominate the polarization signal. As described in Schnittman & Krolik (2009), this transition occurs at lower energies for lower inclinations, as the direct disk contribution is inherently less polarized, so easier to “overcome” by the scattered disk contribution.
@@ 4.2 Prospects for Detection
In this work, we have shown that various QPO mechanisms, including global precession of a tilted torus, orbital modulation by a non-axisymmetric mode, and vertical oscillations in a tilted, truncated disk, can all produce notable changes in polarization. Our hope is that these polarization modulations may be detectable with a dedicated X-ray polarimeter, such as IXPE (Weisskopf et al., 2022) or XPoSat (Paul, 2022). However, we must acknowledge that this would be a very challenging measurement to make in practice. The main difficulty has to do with timescales. Even the low-frequency, type-C QPO that we attribute to precession has frequencies in the range of 0.1-10 Hz, while polarization measurements typically require long exposure times (tens or hundreds of kiloseconds for IXPE). This is obviously many orders of magnitude longer than the QPO period.
@@
As a final point, we note that in this work we have analyzed instantaneous snapshots from our simulations, instead of averaging over realistic exposure times. Thus, the variability in our light curves, especially on short timescales, is likely higher than what would be measured by a real instrument.
@@ 5 Conclusion
We have post-processed the untilted and tilted, truncated black hole accretion disk and isolated, tilted torus simulations of Bollimpalli et al. (2024) using the polarization-capable, Monte Carlo radiation transport code Pandurata  (Schnittman & Krolik, 2013). Our results show reasonable agreement with previous work in that, whenever the disk component dominates, we find a roughly thermal spectrum with an emission peak at a few keV, appropriate for a black hole with $a_{*}=0.9$, disk images that appear nearly symmetric, and polarization degrees that are generally low.
@@
For our untilted, truncated disk, we find very little azimuthal or time dependence, as expected, though we do find a strong inclination dependence. For a nearly edge-on observation, the spectrum can be quite hard, with reflection and coronal scattering dominating at all energies. The polarization degree can also be large ($\gtrsim 2.5$%) in such cases.
@@
We would like to thank the anonymous referee for useful comments and suggestions. We gratefully acknowledge support from SC NASA EPSCoR RGP 2024 and NASA under award No 80NSSC24K0900. DAB acknowledges support from IIT-Indore, through a Young Faculty Research Seed Grant (project: ‘INSIGHT’; IITI/YFRSG/2024-25/Phase-VII/02).